A store had a scratch-and-win promotion in which 1000 cards indicated 10% off, 50 cards indicated 15% off, 10 cards indicated 20% off and 1 card indicated 40% off. What is the expected value for a given scratch-and-win card? *


1
Expert's answer
2021-07-13T13:30:26-0400

"E(X)=\frac{1}{1061}(1000*10+50*15+10*20+1*40)=\frac{10990}{1061}\approx10.36\%."